# Wireless Datasets for Aerial Networks

### v1

@@ Abstract
The integration of unmanned aerial vehicles (UAVs) into 5G-Advanced and future 6G networks presents a transformative opportunity for wireless connectivity, enabling agile deployment and improved LoS communications. However, the effective design and optimization of these aerial networks depend critically on high-quality, empirical data. This paper provides a comprehensive survey of publicly available wireless datasets collected from an airborne platform called Aerial Experimentation and Research Platform on Advanced Wireless (AERPAW). We highlight the unique challenges associated with generating reproducible aerial wireless datasets, and review the existing related works in the literature. Subsequently, for each dataset considered, we explain the hardware and software used, present the dataset format, provide representative results, and discuss how these datasets can be used to conduct additional research. The specific aerial wireless datasets presented include raw I/Q samples from a cellular network over different UAV trajectories, spectrum measurements at different altitudes, flying 4G base station (BS), a 5G-NSA Ericsson network, a LoRaWAN network, an radio frequency (RF) sensor network for source localization, wireless propagation data for various scenarios, and comparison of ray tracing and real-world propagation scenarios. References to all datasets and post-processing scripts are provided to enable full reproducibility of the results. Ultimately, we aim to guide the community toward effective dataset utilization for validating propagation models, developing machine learning algorithms, and advancing the next generation of aerial wireless systems.
@@ I Introduction
Recognizing this potential, regulatory and standardization bodies have begun laying the groundwork for the integration of aerial users into cellular networks. For instance, the Federal Communications Commission (FCC) has examined the use of licensed mid-band spectrum for UAV command-and-control (C2) links and emphasized the role of 5G in enabling unmanned aircraft systems (UAS) integration within cellular networks [4, 5]. In parallel, the 3rd Generation Partnership Project (3GPP) has introduced UAV-specific enhancements in Release 15 and subsequent releases to address challenges such as interference mitigation, mobility management, and flight-relevant key performance indicators (KPIs) for aerial user equipments (UEs) [6]. The International Telecommunication Union (ITU) has similarly recognized the potential of UAVs in advancing global broadband connectivity and enhancing disaster response capabilities [7]. In addition to these regulatory and standardization efforts, industry stakeholders, e.g., Ericsson, Samsung, and Qualcomm, have submitted joint contributions to 3GPP RAN4 aimed at evaluating interference scenarios between aerial and terrestrial users in the shared spectrum bands, highlighting the critical role of empirical measurements in guiding standards development and system design (e.g., see [8]).
This shift toward aerial connectivity has underscored the growing need for open, well-documented, and reproducible wireless datasets captured from airborne platforms. Such datasets serve as a critical enabler for validating theoretical models, training machine learning algorithms, and guiding system-level simulations and protocol design. In practice, producing these datasets requires capabilities beyond those of terrestrial campaigns. These include support for custom waveform generation, programmable radios, repeatable and controllable mobility patterns, and precise synchronization of location and time for each measurement sample.
@@
The remainder of this paper is organized as follows. Section II surveys prior aerial wireless data collection efforts and identifies key gaps and limitations in the existing literature. Section III describes the technical, logistical, and regulatory challenges involved in designing and deploying programmable aerial wireless platforms. Sections V through XIV present a comprehensive set of datasets collected using AERPAW equipment and infrastructure, detailing the measurement configurations, data formats, representative results, and application domains. Finally, Section XV concludes the paper with final remarks and discusses future directions.
II Literature Review
@@
Ensuring future progress in aerial connectivity research requires sustained support for measurement platforms that emphasize openness, documentation, and multi-modal sensing capabilities. The community would benefit from initiatives that align with FAIR data principles (Findable, Accessible, Interoperable, and Reusable) [18], enabling standardized benchmarking and accelerating advancements toward 5G-Advanced and 6G network deployments.
@@ III-B Outdoor Radio Infrastructure and Spectrum
Supporting outdoor experiments with drones and wireless technologies requires access to towers equipped with diverse radio systems and to experimental spectrum bands for testing new waveforms and protocols. While some studies and datasets rely on commercial cellular networks, such approaches limit the scope and flexibility of experiments. The AERPAW platform addresses these challenges through the deployment of five towers, two rooftop sites, and one purpose-built light pole to host USRPs and other commercial off-the-shelf (COTS) wireless devices. All fixed nodes are fiber-connected and dedicated exclusively to AERPAW experimentation. Furthermore, AERPAW is designated as one of the four FCC Innovation Zones in the United States, providing access to specific frequency bands for wireless experimentation with drones [19]. Because airborne transmissions from drones can create significant interference on the ground and with incumbent spectrum users, obtaining experimental frequencies is particularly challenging. AERPAW has secured access to the 900 MHz Industrial, Scientific, and Medical (ISM) band, 1.7/2.1 GHz, and 3.3-3.45 GHz bands to support experiments involving USRPs, commercial 4G/5G equipment, and UAVs, and continues to expand its available experimental spectrum.
@@ III-C Programmable Drones
Another important choice for the AERPAW drones is the open-source software stack employed: in the interest of a fully programmable drone, the AERPAW vehicle control software is fully open: the autopilot firmware is ArduPilot [20], and the software is based on the MAVLink open protocol [21]. The ground control station (GCS) used both in development and operations is QGroundControl [22]. This allows for relying on a large base of existing software while developing software that can be reused by other researchers. The software employed by the AERPAW drones allows experimenters both high-level (e.g., preplanned trajectories) as well as low-level (e.g., off-board control) of the AERPAW drones, allowing for highly customized experiments.
@@ III-D Development in Digital Twins
A unique requirement for AERPAW, among all other PAWR platforms, is its digital twin (DT). In particular, the use of autonomous vehicles in AERPAW and the safety requirements for these vehicles make the development of vehicle software in the testbed itself a major challenge. Instead, for all canonical experiments, all experimenters have to develop their experiments in a custom-made DT of the physical testbed. The AERPAW DT is deployed in the AERPAW data-compute store, which can host several hundred instances of the DT (the exact number depends on the complexity of the experiment instantiated). In the DT, all the software of the real testbed is preserved while simulating three main hardware components of the real testbed: the frames of the drones, the USRPs, and the propagation between the USRPs. The virtual USRPs operate at I/Q sample level, thus allowing for the development of realistic channel models, including antenna patterns, MIMO radios, reflections, and Doppler shifts. The drone emulation includes a virtual machine running the same firmware as the drones on the autopilots of the drones in the testbed, resulting in identical responses to commands of both the drones in the testbed and in the DT.
@@ III-E Precise Localization and Roll/Yaw/Pitch Information
TABLE I: Summary of Datasets by Section, Focus, Data Types, Platforms, Frequency Bands, and Prior Usage.
@@
Name Dataset Link Focus of Dataset Data Types Platforms Frequency Bands Published Papers
@@
Wireless I/Q Datasets (Section V) [24, 25] LTE I/Q sample collection for A2G propagation, channel estimation, spectrum occupancy, and UAV localization SigMF raw I/Q samples; GPS logs; antenna pattern data UAV with NI USRP B205mini SDR; fixed AERPAW radio node (LTE eNB) LTE band [26, 27, 28]
@@
Wireless Spectrum Datasets (Section VI) [29, 30, 31, 32, 33] Wideband aerial spectrum monitoring for propagation analysis, model calibration, and spectrum allocation Power measurements (dBm) with frequency tags; GPS logs; spectrum monitoring logs Helikite with dual NI USRP B205mini-i and dual receive antennas; Intel NUC 87 MHz–6 GHz [34, 35, 36]
@@
5G NSA Wireless KPI Datasets (Section VII) [37, 38, 39, 40] Aerial LTE/NR KPI measurement on a 5G-NSA network using UAV-mounted portable nodes KPI logs (RSRP, RSRQ, SINR, CQI, RI, MCS, throughput); GPS and UAV telemetry AERPAW SAM UAV carrying Quectel 5G modem; Nemo or PawPrints Android device; Ericsson 5G RAN (n77) 5G NR (n77) [41]
@@
LoRa Propagation Datasets (Section VIII) [42] Aerial and ground LoRaWAN propagation for IoT coverage, signal quality, and latency analysis SigMF signal summaries (RSSI, SNR); transmitter and gateway logs (packet metadata, GPS/IMU, spreading factor, data rate); latency metrics Tethered Helikite, UAV, and ground vehicles with USB-programmable LoRa devices; seven LoRaWAN gateways 915 MHz LoRaWAN
@@
Multipath Propagation Datasets (Section IX) [43] A2G multipath channel characterization via synchronized channel sounding (delay spread, Doppler, path loss) Raw I/Q with metadata (GPS, waveform parameters, sync); post-processed CIR and path-loss traces; notebooks and scripts UAV and fixed node with USRP B210; RF front-end; GNSS-disciplined oscillator; Intel NUC 2.4 GHz [44]
@@
Wireless Localization Datasets (TDOA-Based) (Section X) [45, 46] UAV localization and tracking using TDOA from distributed RF sensors under mixed LoS/NLoS conditions; CRLB benchmarking CSV logs of TDOA estimates; GPS ground truth; LoS/NLoS flags; error metrics; UAV telemetry Four Keysight N6841A RF sensors on AERPAW towers; UAV-mounted SDR transmitter or DJI Inspire 2 2.4–3.3 GHz [47, 48]
UAV-Collected RF Measurements for RF Source Localization (AFAR Challenge) (Section XI) [49] UAV and UGV RF localization challenge across DT and real-world testbed for locating an RF-emitting UGV Timestamped RSS and RSQ; synchronized UAV navigation and orientation; CIR-derived quality metrics; experiment metadata UAV and UGV with USRP B205mini SDRs; Intel NUC; DT (V-USRP, CHEM-VM, SITL); AERPAW testbed Cellular spectrum [50], [51, 52, 53]
@@
UAV Signal Classification Datasets (Section XII) [54] RF-based detection and classification of UAV remote controller signals for EW and SIGINT Raw RF captures (MAT files); time-domain samples; spectrograms; signal metadata; MATLAB scripts for preprocessing and database creation Keysight MSOS604A oscilloscope; 24 dBi parabolic antenna; LNA; bandpass filter; MATLAB-based data acquisition system 2.4 GHz ISM band [55, 56, 57]
@@
UAV Trajectory, RSRP, and Throughput Dataset (Section XIII) [58, 59] Trajectory-aware KPI emulation and validation combining emulation and simulation for LTE SISO Processed CSVs (time-stamped lat/lon/alt); RSRP; SNR; throughput; UAV telemetry UAV with LTE SISO radio; AERPAW DT and MATLAB simulation; four BSs LTE SISO [60]
@@
Ray Tracing Simulation and Measurement Comparison Dataset (Section XIV) [61] Direct comparison of measured versus Sionna RT-simulated RSS along UAV trajectories using OSM and Blender environments Measured and simulated RSS; path coefficients; delays; MATLAB alignment scripts UAV transmitter; fixed receivers; NVIDIA Sionna RT 3.3 GHz [62]
Color code: Signal-Level datasets (raw I/Q), Power-Level datasets (spectrum/power measurements), KPI-Level datasets (4G/5G performance metrics).
IV Aerial Wireless Datasets
In the rest of this paper, we will introduce various aerial wireless datasets collected at the NSF AERPAW platform, enabled by the features and capabilities discussed in Section III. For each dataset, we will provide a description of the hardware and software used to collect the dataset, the dataset format, and representative results. References are provided to access each dataset and associated post-processing scripts to replicate the results. Example uses of the dataset are also described.
Table I provides a unified summary of the aerial wireless datasets presented throughout the rest of this article, categorized by their focus, data types, experimental platforms, and prior usage. The color-coded classification, which includes signal-level (I/Q), power-level (RSS), and KPI-level (4G/5G performance) datasets, highlights the breadth and granularity of the measurements.
V Wireless I/Q Datasets

### v2

# Curated Wireless Datasets for Aerial Network Research

@@ Abstract
This Review consolidates publicly available aerial wireless measurement datasets collected using AERPAW. We organize signal-level, power-level, and KPI-level datasets under a unified taxonomy, harmonize metadata, and provide verified access with reproducible post-processing scripts. The curated catalog supports propagation modeling, machine learning, localization, and system-level evaluation for 5G-Advanced and emerging 6G aerial networks.
@@ I Introduction
Recognizing this potential, regulatory and standardization bodies have begun laying the groundwork for the integration of aerial users into cellular networks. For instance, the Federal Communications Commission (FCC) has examined the use of licensed mid-band spectrum for UAV command-and-control (C2) links and emphasized the role of 5G in enabling unmanned aircraft systems (UAS) integration within cellular networks [27, 28]. In parallel, the 3rd Generation Partnership Project (3GPP) has introduced UAV-specific enhancements in Release 15 and subsequent releases to address challenges such as interference mitigation, mobility management, and flight-relevant key performance indicators (KPIs) for aerial user equipments (UEs) [1]. The International Telecommunication Union (ITU) has similarly recognized the potential of UAVs in advancing global broadband connectivity and enhancing disaster response capabilities [40]. In addition to these regulatory and standardization efforts, industry stakeholders, e.g., Ericsson, Samsung, and Qualcomm, have submitted joint contributions to 3GPP Radio Access Network (RAN) Working Group 4 (RAN4) aimed at evaluating interference scenarios between aerial and terrestrial users in the shared spectrum bands, highlighting the critical role of empirical measurements in guiding standards development and system design (e.g., see [20]).
This shift toward aerial connectivity has underscored the growing need for open, well-documented, and reproducible wireless datasets captured from airborne platforms. Such datasets serve as a critical enabler for validating theoretical models, training machine learning algorithms, and guiding system-level simulations, protocol design, and deployment studies. Recent airspace-integration efforts further reinforce this need. For example, the Federal Aviation Administration (FAA) electronic conspicuity white paper outlines near-term pathways that leverage existing terrestrial cellular infrastructure to support UAS functions, such as command-and-control and electronic conspicuity, and highlights multiple open research questions that require operationally representative testing rather than purely simulation-based evaluation [68]. Consequently, curated datasets collected in real environments are increasingly necessary to validate propagation assumptions, quantify KPI behavior at altitude, and inform deployment-relevant designs.
Moreover, the availability of publicly releasable aerial wireless datasets remains limited due to the practical challenges of conducting controlled airborne measurements, including wide-area interference considerations, spectrum coordination requirements, and the need for experimental authorizations. These constraints make it difficult to perform repeatable, large-scale measurement campaigns and highlight the importance of dedicated experimental platforms designed to safely and systematically support real-world aerial data collection.
@@
It is emphasized that this manuscript is intended as a dataset descriptor paper. The contribution does not lie in reporting new experimental findings or previously unobserved channel behaviors. Instead, the novelty resides in the systematic integration, harmonization, and documentation of aerial wireless datasets that have already been collected and publicly released. The paper consolidates these datasets under a unified structure, standardizes their metadata and access mechanisms, and provides reproducible post-processing scripts to facilitate reuse by the research community. As such, the primary objective is to lower the barrier to entry for aerial wireless research by making diverse datasets more findable, accessible, interoperable, and reusable.
The remainder of this paper is organized as follows. The literature review and contributions discuss prior aerial wireless datasets and measurement efforts and identify key gaps motivating this work. The challenges for generating datasets with aerial wireless systems summarize the technical, logistical, and regulatory challenges of collecting reproducible aerial wireless data. The integrated dataset catalog and access presents a unified dataset taxonomy and a consolidated catalog with verified access information. The dataset description sections present the curated AERPAW datasets in a standardized manner, highlighting measurement scope, representative trends, and supported use cases, with detailed file-level information provided in Appendix A. The synthesis, impact, and research outlook consolidates cross-dataset insights, demonstrated research impact, and future research directions. Finally, the paper concludes with concluding remarks.
II Literature Review and Contributions
@@
Ensuring future progress in aerial connectivity research requires sustained support for measurement platforms that emphasize openness, documentation, and multi-modal sensing capabilities. The community would benefit from initiatives that align with FAIR data principles (Findable, Accessible, Interoperable, and Reusable) [88], enabling standardized benchmarking and accelerating advancements toward 5G-Advanced and 6G network deployments. In this work, the presented datasets follow these principles by being findable through publicly indexed repositories with persistent citations, accessible through open download links and accompanying documentation, interoperable via widely used data formats and analysis tools such as MATLAB and Python, and reusable through detailed metadata, experiment descriptions, and provided processing and simulation scripts.
In contrast to prior surveys that focus primarily on analytical models or isolated experimental campaigns, this work complements the state of the art by emphasizing empirical data resources themselves. By situating existing datasets within a unified taxonomy and access framework, the manuscript provides a data-centric perspective on the evolution of aerial wireless research, highlighting trends toward cross-layer measurement, heterogeneous platforms, and reproducible experimentation.
II-A Summary of Contributions
The main contributions of this dataset descriptor are summarized as follows:
  • Integrated dataset taxonomy: A unified categorization of otherwise fragmented aerial wireless datasets into signal-level, power-level, and key performance indicator-level data records, enabling consistent interpretation and cross-dataset comparison across heterogeneous measurement campaigns.
  • Metadata harmonization: Standardized and consistent documentation of measurement parameters, including platform altitude, geographic context, carrier frequency, bandwidth, sampling configuration, and calibration status, addressing common interoperability and reproducibility limitations in existing public datasets.
  • Accessible and reproducible data access: Consolidation of publicly available datasets into a single reference with verified access links or persistent repositories, accompanied by open-source post-processing and visualization scripts to support reproducible and transparent reuse.
  • Validation and usage guidance: Representative data visualizations and analysis examples provided as integrity checks and practical guidance, clarifying intended use cases and reuse potential for propagation modeling, system evaluation, localization, and data-driven research.
  • FAIR-oriented dataset design: Alignment with the principles of findability, accessibility, interoperability, and reusability to facilitate long-term benchmarking and comparative research within the aerial wireless community.
@@ III-B Outdoor Radio Infrastructure and Spectrum
Supporting outdoor experiments with drones and wireless technologies requires access to towers equipped with diverse radio systems and to experimental spectrum bands for testing new waveforms and protocols. While some studies and datasets rely on commercial cellular networks, such approaches limit the scope and flexibility of experiments. The AERPAW platform addresses these challenges through the deployment of five towers, two rooftop sites, and one purpose-built light pole to host USRPs and other COTS wireless devices. All fixed nodes are fiber-connected and dedicated exclusively to AERPAW experimentation. Furthermore, AERPAW is designated as one of the four FCC Innovation Zones in the United States, providing access to specific frequency bands for wireless experimentation with drones [6]. Because airborne transmissions from drones can create significant interference on the ground and with incumbent spectrum users, obtaining experimental frequencies is particularly challenging. AERPAW has secured access to the 900 MHz Industrial, Scientific, and Medical (ISM) band, 1.7/2.1 GHz, and 3.3-3.45 GHz bands to support experiments involving USRPs, commercial 4G/5G equipment, and UAVs, and continues to expand its available experimental spectrum.
@@ III-C Programmable Drones
Another important choice for the AERPAW drones is the open-source software stack employed: in the interest of a fully programmable drone, the AERPAW vehicle control software is fully open: the autopilot firmware is ArduPilot [ArduPilot], and the software is based on the MAVLink open protocol [MAVLink]. The ground control station (GCS) used both in development and operations is QGroundControl [QGroundControl]. This allows for relying on a large base of existing software while developing software that can be reused by other researchers. The software employed by the AERPAW drones allows experimenters both high-level (e.g., preplanned trajectories) as well as low-level (e.g., off-board control) of the AERPAW drones, allowing for highly customized experiments.
@@ III-D Development in Digital Twins
A unique requirement for AERPAW, among all other PAWR platforms, is its digital twin (DT). In particular, the use of autonomous vehicles in AERPAW and the safety requirements for these vehicles make the development of vehicle software in the testbed itself a major challenge. Instead, for all canonical experiments, all experimenters have to develop their experiments in a custom-made DT of the physical testbed. The AERPAW DT is deployed in the AERPAW data-compute store, which can host several hundred instances of the DT (the exact number depends on the complexity of the experiment instantiated). In the DT, all the software of the real testbed is preserved while simulating three main hardware components of the real testbed: the frames of the drones, the USRPs, and the propagation between the USRPs. The virtual USRPs operate at I/Q sample level, thus allowing for the development of realistic channel models, including antenna patterns, Multiple-Input Multiple-Output (MIMO) radios, reflections, and Doppler shifts. The drone emulation includes a virtual machine running the same firmware as the drones on the autopilots of the drones in the testbed, resulting in identical responses to commands of both the drones in the testbed and in the DT.
@@ III-E Precise Localization and Roll/Yaw/Pitch Information
TABLE I: Unified and verified summary of publicly available aerial wireless datasets [3], including persistent DOIs or access links, data formats, platforms, frequency bands, and representative prior usage to support reproducible research.
@@
Name DOI / Access Link Focus of Dataset Data Types and Formats Platforms Frequency Published Papers
@@
Wireless I/Q Datasets (Signal-Level) [52, 54] LTE I/Q sample collection for A2G propagation, channel estimation, spectrum occupancy, and UAV localization SigMF raw I/Q samples; GPS logs; antenna pattern metadata UAV with NI USRP B205mini; fixed AERPAW LTE eNB LTE [60, 55, 51]
@@
Wireless Spectrum Datasets (Power-Level) [76, 75, 57, 59, 58, 77] Wideband aerial spectrum monitoring for propagation analysis and model calibration PSD (dBm); frequency-tagged logs; GPS metadata Helikite with dual USRP B205mini-i; Intel NUC 87 MHz–6 GHz [73, 74]
@@
5G NSA Wireless KPI Datasets (KPI-Level) [10, 83, 82] Aerial LTE/NR KPI measurements on a 5G-NSA network using UAV-mounted nodes KPI logs (RSRP, RSRQ, SINR, throughput); GPS and UAV telemetry (CSV/JSON) AERPAW SAM UAV; Quectel 5G modem; Ericsson 5G RAN n77 [11]
@@
LoRa Propagation Datasets (Power-Level) [87] Aerial and ground LoRaWAN propagation for IoT coverage and latency analysis RSSI/SNR; packet metadata; latency metrics; GPS/IMU logs UAV, tethered helikite, ground vehicles; LoRaWAN gateways 915 MHz
@@
Multipath Propagation Datasets (Signal-Level) [35] A2G multipath channel characterization via synchronized channel sounding Raw I/Q with metadata; CIRs; path-loss traces; notebooks UAV and fixed node with USRP B210; GNSS-disciplined oscillators 2.4 GHz [34]
@@
Wireless Localization Datasets (TDOA-Based) (Signal-Level) [18] UAV localization and tracking using distributed RF sensors TDOA estimates; GPS ground truth; LoS/NLoS labels; error metrics Keysight RF sensors; UAV-mounted SDR 2.4–3.3 GHz [17]
AFAR RF Source Localization Dataset (Power-Level) [33] UAV/UGV RF source localization across digital twin and testbed environments RSS/RSQ; UAV navigation; CIR-derived metrics; metadata UAV and UGV with USRP B205mini; AERPAW digital twin and testbed Cellular [61, 46, 62]
@@
UAV Signal Classification Datasets (Signal-Level) [22] RF-based detection and classification of UAV controller signals RF captures (MAT); spectrograms; labels; MATLAB scripts Keysight oscilloscope; parabolic antenna; RF front-end 2.4 GHz [23, 24]
@@
UAV Trajectory, RSRP, and Throughput Dataset (KPI-Level) [86, 38] Trajectory-aware KPI emulation and validation for LTE SISO links Time-stamped CSV; RSRP; SNR; throughput; UAV telemetry UAV LTE SISO; AERPAW digital twin; MATLAB simulation LTE [37]
@@
Ray-Tracing Measurement Comparison Dataset (Power-Level) [84] Comparison of measured and ray-tracing-simulated RSS along UAV trajectories Measured/simulated RSS; path coefficients; delays; MATLAB scripts UAV transmitter; fixed receivers; NVIDIA Sionna RT 3.3 GHz [50]
IV Integrated Dataset Catalog and Access
This work consolidates multiple aerial wireless datasets that were collected across different campaigns, platforms, and experimental objectives, as summarized in Table I. To facilitate integration and reuse, all datasets are organized under a unified taxonomy based on the level of abstraction at which the measurements are recorded.
At the lowest level, signal-level datasets provide raw in-phase and quadrature samples suitable for detailed physical-layer analysis and custom receiver processing. Power-level datasets capture received signal strength and path loss behavior as functions of altitude, distance, and environment. Key performance indicator-level datasets contain higher-layer metrics such as throughput, latency, and packet error rate, enabling system-level evaluation.
Each dataset is accompanied by a standardized metadata record that includes measurement location, altitude range, carrier frequency, bandwidth, antenna configuration, and temporal resolution. Where applicable, data are provided in widely adopted formats to promote interoperability, and conversion scripts are supplied when proprietary or campaign-specific formats are used.
All datasets and associated scripts are hosted in publicly accessible repositories with persistent links. This centralized catalog structure enables users to efficiently identify relevant datasets, understand their characteristics, and integrate them into their own analysis pipelines.
To facilitate informed dataset selection and comparative evaluation, Table II provides a task-oriented comparison of the surveyed datasets. This comparison highlights practical considerations that emerge from the analysis, including the importance of matching dataset abstraction level to the intended research objective, recognizing trade-offs between raw signal fidelity and system-level scalability, and accounting for dataset-specific limitations such as spatial coverage, technology dependence, and environmental specificity when interpreting results or combining datasets. For instance, we explicitly recommend utilizing Signal-Level datasets for physical layer algorithm development where phase information is critical, while advising the use of KPI-Level datasets for scalable network orchestration studies where storage efficiency and processing speed are paramount.
Building on this unified catalog and comparative perspective, we next turn to the individual datasets in detail. The following sections present a set of aerial wireless datasets collected at the NSF AERPAW, enabled by the experimental capabilities and constraints described earlier. For each dataset, we summarize the measurement objectives, hardware and software configuration, dataset format, and representative results. Public repository links and post-processing scripts are provided to support reproducibility and reuse, and selected application examples are included to illustrate relevance to propagation analysis, system evaluation, and data-driven aerial wireless research.
TABLE II: Task-oriented comparative analysis of aerial wireless datasets, with dataset categories annotated by manuscript section labels to support cross-dataset selection and reuse.
[…]
Legend:  well suited,  $\triangle$ partially suited,  $\times$ not suited.
V Wireless I/Q Dataset